OpenStudy (anonymous):

The percent composition of all elements in HClO3 is: a. 2% H, 20% Cl, 78% O b. 1% H, 42% Cl, 57% O c. 2% H, 68% Cl, 30% O d. None of the above

OpenStudy (anonymous):

well take molar mass of HClO3 and atomic masses of atoms in compound and divide each atomic mass with molar mass of compound

OpenStudy (anonymous):

idk how to do that, im completelylost because it has different elements ombined.

OpenStudy (anonymous):

H - 1,008 Cl - 35,45 O - 15,999 M(HClO3) = (3* 15,999) + 1,008 + 35,45 = 84,455 x(H) = 1,008/84,455 =0,0119 = 1,2 % x(Cl) = 35,45/ 84,455= 0,419 = 42% x(O)= 3*15,999 / 84,455=0,5683 = 57 %

OpenStudy (anonymous):

so answer is: b. 1% H, 42% Cl, 57% O

OpenStudy (anonymous):

thank you so muchchemistry is so confusing. im still in the beginning of this section with mole calculations and i still dont understand it.

OpenStudy (anonymous):

read it few times so it settles down in your head, after you learn basics chemistry is pretty easy... :)